Two numbera that's difference is 3 but its sum is 45
Found 2 solutions by jim_thompson5910, oberobic:
Answer by jim_thompson5910(28536)

x-y=3
x+y=45

Add the two equations to get

2x=48

Now solve for x

2x=48
x=48/2
x=24

Now use this to find y

x-y=3
24-y=3
-y=3-24
-y=-21
y=21

So the solutions are x=24 and y=21

So the two numbers are 24 and 21

If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

Answer by oberobic(2304)

x + y = 45
y = x+3
.
substitute
.
x +(x+3) = 45
2x = 42
x = 21
y = x+3 = 24
.
The two numbers are 21 and 24.
.
Done.